OpenStudy (anonymous):

the width is X the height is Y the question says find the volume(v) of the box as a function of X

jhonyy9 (jhonyy9):

hope the base of this box is a square so than the volume will be area of base time height so V= x*x*y = x^2 *y hope this is understandably